OpenStudy (anonymous):

What is the 46th term of the arithmetic series 1, 8, 15, … ? 181 243 316 422

OpenStudy (help!!!!):

looks like 7 at first but it not.......

OpenStudy (anonymous):

Tn= a+(n-1)d where Tn is the term , a = 1st term of Arithmetic progression d is common difference

OpenStudy (anonymous):

so T46 = 1+(46-1)*7= 1+ (45*7) = 316

OpenStudy (anonymous):

\[a_{n} = 1 + (n-1) \cdot 7 \implies a_{46} = 1 + (46-1)\cdot 7 = 316\]

OpenStudy (anonymous):

Aw thanks guys! :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):